# Variational Bayesian Optimal Experimental Design with Normalizing Flows

### v1

@@ 1 Introduction
Mathematical models are central for understanding and making predictions in physical systems. Parameters in models are often uncertain, but their uncertainty can be reduced when information is gained from experimental observations. Since experiments are costly in many applications, optimal experimental design (OED) (see, e.g., [1, 2, 3, 4, 5]) can bring substantial resource savings by identifying the experiments that produce the most valuable data.
@@
Bayesian OED further incorporates the prior and posterior uncertainty in its objective function. The seminal paper by Lindley [6] proposed to use mutual information (MI) between model parameters and experimental data as the design criterion, which is equivalent to the expected information gain (EIG) in the model parameters. The EIG, however, is generally intractable to compute and has to be approximated numerically. One strategy entails directly approximating the EIG. For example, Ryan [7] introduced a nested Monte Carlo (NMC) estimator for the EIG, but it is computationally expensive with the number of likelihood evaluations scaling as the product of sample sizes of the nested Monte Carlo loops. Advanced numerics have been developed to accelerate NMC computations through surrogate modeling [8, 9], reusing samples across nested loops [8], and ‘Gaussianizing’ the posterior via Laplace approximations [10, 11, 12, 13]. Another estimator similar to the NMC is the prior contrastive estimator (PCE) [14]. A key difference is that PCE reuses each outer loop sample once in the inner loop, resulting the estimator mean to be negatively biased (i.e., a lower bound of the EIG). The above techniques, however, all require the ability to evaluate the likelihood function, and cannot accommodate implicit or intractable likelihood settings that often arise from stochastic models. An EIG lower bound that is able to handle intractable likelihood is the LB-KLD [15], which can be derived from Shannon’s entropy power inequality.
@@
Another strategy is to adopt a variational bound for the EIG [16], and tighten the bound by optimizing its variational parameters. (Note that PCE and LBKLD do not have variational parameters for tightening the bounds, and therefore are not considered to be variational bounds.) We call these variational OED (vOED) approaches. For example, the tractable unnormalized Barber–Agakov (TUBA) lower bound [16] incorporates the tuning of a ‘critic’ function, and the Nguyen–Wainwright–Jordan (NWJ) bound [17], also known as the mutual information neural estimation f-divergence (MINE-f) bound [18], can be shown to be a special case of TUBA and has been used by Kleinegesse and Gutmann [19] in the OED context. Another lower bound, the information noise-contrastive estimation (InfoNCE) bound [20], takes a similar form as PCE, but replaces the likelihood with an exponentiated critic function. More recently, Ivanova et al. [21] extended the NWJ and InfoNCE bounds to the sequential OED setting. All of these bounds can accommodate implicit likelihood.
@@
A particular type of variational bounds, known as the Barber–Agakov (BA) lower bound [22], emerges from approximating the posterior density function directly. Foster et al. [23] were the first to use the BA bound in the OED setting, but remained with relatively simple variational distributions such as Gaussian, Bernoulli, and Gamma. The approximations thus can deteriorate when the true posterior distributions depart from these forms. Improving upon their work, Foster et al. [14] then introduced the adaptive contrastive estimation (ACE) lower bound based on the PCE but used a variational biasing distribution for the inner loop sampling. However, ACE required explicit likelihood.
@@
One of the key contributions of this paper is therefore to reduce the BA bound bias in estimating the EIG by improving the accuracy of posterior approximation through enriching the space of variational distributions using normalizing flows (NFs) [24, 25, 26, 27]. We achieve this while maintaining the ability to handle implicit likelihood.
@@
Different structures and parameterizations of transport maps have been proposed. Triangular and block-triangular forms [32], often parameterized using multivariate polynomials [33, 29], radial basis functions [29], tensor decomposition [34], and partially convex potential maps and conditional optimal transport flows [35]. Invertibility is then enforced through local monotonicity constraints, or setting parameterizations in ways to guarantee monotonicity [29, 35]. For instance, Baptista et al. [36] proposed a rectification operator that can always transform sufficiently smooth non-monotone functions into monotone component functions of a triangular map, and also demonstrated the elimination of spurious local minima that often arises in the training of neural network-based maps. Furthermore, the triangular structures provide a convenient means to extract conditional densities (e.g., the posterior or likelihood) and to compute the Jacobian of transformation.
Elsewhere, efforts in the machine learning community in creating practical transport maps focused on composing together invertible, often neural network-based, functions, known as normalizing flows. For recent reviews of NFs, see [27, 26, 26]. While first represented as a simple 3-trainable-parameter function [37] and then as a composition of localized radial expansion to achieve greater expressiveness [38], the architectural form of NFs has evolved rapidly. Examples include planar and radial flows [25, 39, 40, 41], coupling layers and autoregressive flows [42, 24, 43, 44, 45], splines [46], residual flows [47] and neural ordinary differential equations [48]. While NFs have been used for approximating the posterior in a variational inference context ([25] and many subsequent works), they have not yet been explored in OED.
@@ 2.2 Variational optimal experimental design
When $U_{L}(\mathbf{d};\boldsymbol{\lambda})$ is differentiate with respect to $\mathbf{d}$ and $\boldsymbol{\lambda}$, their gradients are
@@ 3 Computational Methods
where samples $\boldsymbol{\theta}^{(i)}\sim p(\boldsymbol{\theta})$, $\mathbf{y}^{(i)}\sim p(\mathbf{y}|\boldsymbol{\theta}^{(i)},\mathbf{d})$, and $\boldsymbol{\theta}^{(i,j)}\sim p(\boldsymbol{\theta})$ are drawn from the prior and likelihood. Calculating the NMC estimator $\widehat{U}_{\text{NMC}}$ requires evaluating the likelihood $p(\mathbf{y}|\boldsymbol{\theta},\mathbf{d})$ and cannot directly handle implicit likelihood settings. $\widehat{U}_{\text{NMC}}$ is a biased estimator to $U$ under any finite inner-loop sample size $N_{\text{in}}$, but it is asymptotically unbiased as $N_{\text{in}}\rightarrow\infty$. Properties of the NMC estimator has been studied extensively [7, 49, 13, 50], and we will use it for reference comparison in this work when applicable.
For vOED, the lower and upper bounds in Eqn. 5 and 13 require numerical approximations for their expectation operators, for example through standard MC estimation:
@@
where $\boldsymbol{\theta}^{(i)}\sim p(\boldsymbol{\theta})$ and $\mathbf{y}^{(i)}\sim p(\mathbf{y}|\boldsymbol{\theta}^{(i)},\mathbf{d})$. In particular, calculating Eqn. 19 only requires sampling the likelihood without any likelihood evaluations; it is therefore suitable to handle implicit likelihood cases. Note that both $\widehat{U}_{L}$ and $\widehat{U}_{U}$ are not bounds themselves, but are (unbiased) estimators of the bounds $U_{L}$ and $U_{U}$, respectively. With respect to the true EIG $U$, then, $\widehat{U}_{L}$ and $\widehat{U}_{U}$ are biased estimators whose bias are, respectively, negative and positive.
@@ 3.1 Normalizing flows
$f_{2}(f_{1}(\widetilde{\boldsymbol{\Theta}}))$ $=\begin{bmatrix}\mathbf{Z}_{1}=\widetilde{\boldsymbol{\Theta}}_{1}\odot\exp(s_{2}(\mathbf{Z}_{2}))+t_{2}(\mathbf{Z}_{2})\\ {\boldsymbol{\Theta}^{\prime}}_{2}\end{bmatrix}.$ (28)
@@ 4 Numerical Experiments
We demonstrate vOED-NFs across four design cases involving models of varying complexity: (1) a nonlinear model with 3 parameters and 1 design variable; (2) a linear model with 21 parameters and 400-dimensional design vector; (3) a PDE model for cathodic electrophoretic deposition; and (4) a stochastic model of aphid population with implicit likelihood.
@@ 4.1 Case 1: Low-dimensional nonlinear design
Consider a nonlinear observation model
$y=G(\boldsymbol{\theta},d)+\epsilon\qquad\text{where}\qquad G(\boldsymbol{\theta},d)=\theta_{1}^{3}d^{2}+\theta_{2}e^{-|0.2-d|}+\sqrt{\theta_{3}^{2}d^{2}},$ (30)
@@
Figure 2(a) presents the EIG estimates across the discretized design space using (a) (Hi-NMC) a high-quality $\widehat{U}_{\text{NMC}}$ with $N_{\text{out}}=20000$ and $N_{\text{in}}=20000$ samples, (b) (Lo-NMC) a low-quality $\widehat{U}_{\text{NMC}}$ with $N_{\text{out}}=200$ and $N_{\text{in}}=200$ samples, (c) (vOED-G) EIG lower bound $\widehat{U}_{L}$ using Gaussian distributions optimized with $N_{\text{opt}}=20000$ samples, and (d) (vOED-NFs) EIG lower bound $\widehat{U}_{L}$ using NFs optimized with $N_{\text{opt}}=20000$ samples. Furthermore, the lower bound estimate $\widehat{U}_{L}$ at each $d$ is evaluated using $N=10000$ samples. vOED-NFs adopts $T=5$ sets of complete transformations of the cINN described in Sec. 3.1. Additional training details and hyperparameter choices for (c) and (d) can be found in Sec. B.1.
In Fig. 2(a), Hi-NMC is designated as the reference solution. In comparison to this reference, Lo-NMC exhibits moderate bias while vOED-NFs performs extremely well. vOED-G, due to its much simplistic Gaussian variational distributions, deviates significantly from the expected utility trend and misidentifies the optimal design. However, the training cost for vOED-NFs is also higher than vOED-G, as illustrated through the training convergence history for $d=1.0$ in Fig. 2(b). Here vOED-G shows a very rapid convergence to $\boldsymbol{\lambda}^{*}$, while vOED-NFs takes longer to stabilize and experiences greater fluctuation. This timing difference is negligible when the overall computations are dominated by the forward model runs, but can become significant when the forward model is inexpensive.
Figure 2: Case 1. EIG estimates and sample training convergence plot.
@@
Figure 5 investigates the impact of the number of transformations and optimization sample size $N_{\text{opt}}$ in vOED-NFs. The plots show the EIG lower bound estimates $\widehat{U}_{L}$ when adopting $T=\{1,3,5,7\}$ sets of complete cINN transformations, and using $N_{\text{opt}}=\{5000,10000,50000\}$. The EIG lower bound estimates are still evaluated with $N=10000$ samples. In general, employing 3 transformations is adequate to identify the optimal design $d^{\ast}=1$. However, when $N_{\text{opt}}$ is small, adding more transformations does not provide significant improvement. When $N_{\text{opt}}$ is increased, the value from adding transformations is more notable and stabilized around $T=5$; this is consistent with previous work [52] that reported composing more than 5 transformations provides only slight additional benefits. This diminishing return is likely due to the much larger number of NFs parameters when a high number of transformations is used, making them more prone to overfitting. Additional testing and discussions for the $N_{\text{opt}}=5000$ case can be found in B.1.
@@ 4.2 Case 2: High-dimensional linear design
where $y_{j}\in\mathbb{R}$ is the $j$th observation, $\mathbf{d}_{j}\in\mathbb{R}^{p}$ is its corresponding design vector, $\epsilon_{j}\sim\mathscr{N}(0,\sigma^{2})$ is a Gaussian noise random variable, and $\boldsymbol{\theta}=\{\mathbf{w},\sigma\}$ encompasses the $p$-dimensional regression coefficient vector $\mathbf{w}$ and noise standard deviation $\sigma$. We set $p=20$ and $n=20$, hence the parameter dimension is 20 and the total design dimension is 400. Independent priors are adopted for $\mathbf{w}_{j}\sim\text{Laplace}(0,1)$ and $\sigma\sim\text{Exp}(1)$. Lastly, a design constraint $||\mathbf{d}_{j}||_{1}=1$ is imposed to reflect a resource budget; this is implemented through re-normalization of $\mathbf{d}$ after each SGA update in order to map it back to the feasible region. During the joint optimization of $\mathbf{d}$ and $\boldsymbol{\lambda}$, the total number of forward model runs (i.e., the total number of $(\boldsymbol{\theta}^{(i)}$, $\mathbf{y}^{(i)})$ pairs across all design epochs) is fixed to $10^{6}$. This is the same computational budget used by [14] that adopted Gaussian and Gamma variational distributions for $\mathbf{w}$ and $\sigma$, respectively; we refer to the setup in [14] as vOED-GG. Further details on the computational setup can be found in C.1.
Optimization results are presented in Table 1, with the middle column showing the mean $\pm$ one standard deviation of 10 lower bound estimates $\widehat{U}_{L}$ (each evaluated with $N=10000$ samples) at the optimal design found by vOED-GG and vOED-NFs, and the right column showing the mean $\pm$ one standard deviation of 10 high-quality NMC estimates (each using $N_{\text{out}}=10000$ and $N_{\text{in}}=10000$) at each method’s optimal design to provide an accurate EIG estimate and using a common estimation method. We observe that vOED-NFs’s optimal design achieves a tighter lower bound estimate and also a higher NMC estimate, indicating that a better design has been found than vOED-GG. However, the vOED-NFs lower bound estimate is not tight, and the remaining discrepancy between $\widehat{U}_{L}$ and $\widehat{U}_{\text{NMC}}$ may be contributed by a number of factors: (1) the functional form of cINN; (2) the number of cINN transformations; (3) the functional representation of $s$’s and $t$’s; (4) the possibility to converge to a local optimum during optimization; and (5) the finite sample sizes.
@@
vOED-GG   12.07 $\pm$ 0.11 $24.28\pm 0.24$
vOED-NFs   20.94 $\pm$ 0.19 $24.91\pm 0.24$
Table 1: Case 2. Middle column shows the mean of 10 lower bound estimates $\widehat{U}_{L}$ (each is evaluated with $N=10000$ samples) at the optimal design found by vOED-GG and vOED-NFs, and the right column shows the mean of 10 high-quality NMC estimates (each with $N_{\text{out}}=10000$ and $N_{\text{in}}=10000$) at the optimal designs. The $\pm$ values are one standard deviation estimates for the estimators.
In order to compare the posterior approximation capabilities of the two approaches, we look at their approximate posteriors at a common design with each $\mathbf{d}_{j}=\mathbf{e}_{j}$ set to the standard unit vector in the $j$th dimension, and at three $\mathbf{y}$ samples. Figures 6 and 7 present the marginal posteriors for $\{w_{1},w_{2},w_{3},w_{4},\sigma\}$ using No-U-Turn Hamiltonian Monte Carlo (NUTS-HMC) [58] as the reference posterior distribution, along with the approximations obtained from vOED-GG and vOED-NFs. From the figure, it is evident that the vOED-NFs posteriors can capture the non-Gaussian structures much better than vOED-GG’s Gaussian/Gamma distributions.
Figure 6: Case 2. Comparison of marginal posteriors obtained from NUTS-HMC and vOED-GG [23] at design where $\mathbf{d}_{j}=\mathbf{e}_{j}$.
Figure 7: Case 2. Comparison of marginal posteriors obtained from NUTS-HMC and vOED-NFs [23] at design where $\mathbf{d}_{j}=\mathbf{e}_{j}$.

### v2

@@ 1 Introduction
Mathematical models are central for understanding and making predictions in physical systems. Parameters in models are often uncertain, but their uncertainty can be reduced when information is gained from experimental observations. Since experiments are costly in many applications, optimal experimental design (OED) (see, e.g., [1, 2, 3, 4, 5, 6]) can bring substantial resource savings by identifying the experiments that produce the most valuable data.
@@
Bayesian OED further incorporates the prior and posterior uncertainty in its objective function. The seminal paper by Lindley [7] proposed to use mutual information (MI) between model parameters and experimental data as the design criterion, which is equivalent to the expected information gain (EIG) in the model parameters. The EIG, however, is generally intractable to compute and has to be approximated numerically. One strategy entails directly approximating the EIG. For example, Ryan [8] introduced a nested Monte Carlo (NMC) estimator for the EIG, but it is computationally expensive with the number of likelihood evaluations scaling as the product of sample sizes of the nested Monte Carlo loops. Advanced numerics have been developed to accelerate NMC computations through surrogate modeling [9, 10], reusing samples across nested loops [9], and ‘Gaussianizing’ the posterior via Laplace approximations [11, 12, 13, 14]. Another estimator similar to the NMC is the prior contrastive estimator (PCE) [15]. A key difference is that PCE reuses each outer loop sample once in the inner loop, resulting the estimator mean to be negatively biased (i.e., a lower bound of the EIG). The above techniques, however, all require the ability to evaluate the likelihood function and cannot accommodate implicit or intractable likelihood settings that often arise from stochastic models. An EIG lower bound that is able to handle intractable likelihood is the LB-KLD [16], which can be derived from Shannon’s entropy power inequality.
@@
Another strategy is to adopt a variational bound for the EIG [17] and tighten the bound by optimizing its variational parameters. (Note that PCE and LBKLD do not have variational parameters for tightening the bounds, and therefore are not considered to be variational bounds.) We call these variational OED (vOED) approaches. For example, the tractable unnormalized Barber–Agakov (TUBA) lower bound [17] incorporates the tuning of a ‘critic’ function, and the Nguyen–Wainwright–Jordan (NWJ) bound [18], also known as the mutual information neural estimation f-divergence (MINE-f) bound [19], can be shown to be a special case of TUBA and has been used by Kleinegesse and Gutmann [20] in the OED context. Another lower bound, the information noise-contrastive estimation (InfoNCE) bound [21], takes a similar form as PCE, but replaces the likelihood with an exponentiated critic function. More recently, Ivanova et al. [22] extended the NWJ and InfoNCE bounds to the sequential OED setting. All of these bounds can accommodate implicit likelihoods.
@@
A particular type of variational bounds, known as the Barber–Agakov (BA) lower bound [23], emerges from approximating the posterior density function directly. A potential advantage of the BA bound is its ability to enable both density evaluation and sampling from the resulting approximate posteriors. This is in contrast to several other lower bound choices, such as the NWJ bound that allows density evaluation but not sampling, and the InfoNCE bound that does not offer density or sampling. Foster et al. [24] were the first to use the BA bound in the OED setting, but remained with relatively simple variational distributions such as Gaussian, Bernoulli, and Gamma. The approximations thus can deteriorate when the true posterior distributions depart from these forms. Improving upon their work, Foster et al. [15] then introduced the adaptive contrastive estimation (ACE) lower bound based on the PCE but used a variational biasing distribution for the inner loop sampling. However, ACE required explicit likelihood.
@@
One of the key contributions of this paper is therefore to reduce the BA bound bias in estimating the EIG by improving the accuracy of posterior approximation through enriching the space of variational distributions using normalizing flows (NFs) [25, 26, 27, 28]. We achieve this while maintaining the ability to handle implicit likelihoods.
@@
Different structures and parameterizations of transport maps have been proposed. Triangular and block-triangular forms [33] are often parameterized using multivariate polynomials [34, 30], radial basis functions [30], tensor decomposition [35], and partially convex potential maps and conditional optimal transport flows [36]. Invertibility is then enforced through local monotonicity constraints, or setting parameterizations in ways to guarantee monotonicity [30, 36]. For instance, Baptista et al. [37] proposed a rectification operator that can always transform sufficiently smooth non-monotone functions into monotone component functions of a triangular map, and also demonstrated the elimination of spurious local minima that often arises in the training of neural network-based maps. Furthermore, the triangular structures provide a convenient means to extract conditional densities (e.g., the posterior or likelihood) and to compute the Jacobian of transformation.
Elsewhere, efforts in the machine learning community in creating practical transport maps focused on composing together invertible, often neural network-based, functions, known as normalizing flows. For recent reviews of NFs, see [28, 27, 27]. While first represented as a simple 3-trainable-parameter function [38] and then as a composition of localized radial expansion to achieve greater expressiveness [39], the architectural form of NFs has evolved rapidly. Examples include planar and radial flows [26, 40, 41, 42], coupling layers and autoregressive flows [43, 25, 44, 45, 46], splines [47], residual flows [48] and neural ordinary differential equations [49]. While NFs have been used for approximating the posterior in a variational inference context (see [26] and several subsequent works), their application in OED have been only lightly explored. Kennamer et al. [50] demonstrated the use of NFs for EIG estimation focusing on generalized linear models (GLMs), while Orozco et al. [51] applied NFs to OED for image datasets and studied posterior samples of the test error—neither has yet made comparisons with alternative OED estimators, such as NMC.
@@ 2.2 Variational optimal experimental design
When $U_{L}(\mathbf{d};\boldsymbol{\lambda})$ is differentiated with respect to $\mathbf{d}$ and $\boldsymbol{\lambda}$, their gradients are
@@ 3 Computational Methods
where samples $\boldsymbol{\theta}^{(i)}\sim p(\boldsymbol{\theta})$, $\mathbf{y}^{(i)}\sim p(\mathbf{y}|\boldsymbol{\theta}^{(i)},\mathbf{d})$, and $\boldsymbol{\theta}^{(i,j)}\sim p(\boldsymbol{\theta})$ are drawn from the prior and likelihood. Calculating the NMC estimator $\widehat{U}_{\text{NMC}}$ requires evaluating the likelihood $p(\mathbf{y}|\boldsymbol{\theta},\mathbf{d})$ and cannot directly handle implicit likelihoods settings. $\widehat{U}_{\text{NMC}}$ is a biased estimator to $U$ under any finite inner-loop sample size $N_{\text{in}}$, but it is asymptotically unbiased as $N_{\text{in}}\rightarrow\infty$. Properties of the NMC estimator has been studied extensively [8, 52, 14, 53], and we will use it for reference comparison in this work when applicable.
For vOED, the lower and upper bounds in Eqn. 5 and 13 require numerical approximations for their expectation operators, for example through standard Monte Carlo (MC) estimation:
@@
where $\boldsymbol{\theta}^{(i)}\sim p(\boldsymbol{\theta})$ and $\mathbf{y}^{(i)}\sim p(\mathbf{y}|\boldsymbol{\theta}^{(i)},\mathbf{d})$. In particular, calculating Eqn. 19 only requires sampling the likelihood without any likelihood evaluations; it is therefore suitable to handle implicit likelihoods. Note that both $\widehat{U}_{L}$ and $\widehat{U}_{U}$ are not bounds themselves, but are (unbiased) estimators of the bounds $U_{L}$ and $U_{U}$, respectively. With respect to the true EIG $U$, then, $\widehat{U}_{L}$ and $\widehat{U}_{U}$ are biased estimators whose biases are, respectively, negative and positive.
@@ 3.1 Normalizing flows
$f_{2}(f_{1}(\widetilde{\boldsymbol{\Theta}}))$ $=\begin{bmatrix}\mathbf{Z}_{1}=\widetilde{\boldsymbol{\Theta}}_{1}\odot\exp(s_{2}(\mathbf{Z}_{2}))+t_{2}(\mathbf{Z}_{2})\\ \mathbf{Z}_{2}\end{bmatrix}.$ (28)
@@
In summary, the final approach to determine the optimal design entails maximizing the EIG via its lower bound:
$\mathbf{d}^{\ast},\boldsymbol{\lambda}^{\ast}=\operatornamewithlimits{argmax}_{\mathbf{d}\in\mathscr{D},\boldsymbol{\lambda}}\widehat{U}_{L}=\operatornamewithlimits{argmax}_{\mathbf{d},\boldsymbol{\lambda}}\frac{1}{N}\sum_{i=1}^{N}\left\{\ln q(\boldsymbol{\theta}^{(i)}|\mathbf{y}^{(i)};\boldsymbol{\lambda})-\ln p(\boldsymbol{\theta}^{(i)})\right\},$ (30)
where $\boldsymbol{\theta}^{(i)}\sim p(\boldsymbol{\theta})$, and $\mathbf{y}^{(i)}\sim p(\mathbf{y}^{(i)}|\boldsymbol{\theta}^{(i)},\mathbf{d})$, and $\boldsymbol{\lambda}$ represents all the neural network weight and bias parameters in the cINN and summary network. For results that show upper bound estimates, they are based on solving the problem:
$\mathbf{d}^{\ast}=\operatornamewithlimits{argmax}_{\mathbf{d}\in\mathscr{D}}\min_{\boldsymbol{\lambda}}\widehat{U}_{U}=\operatornamewithlimits{argmax}_{\mathbf{d}\in\mathscr{D}}\min_{\boldsymbol{\lambda}}\frac{1}{N}\sum_{i=1}^{N}\left\{\ln p(\mathbf{y}^{(i)}|\boldsymbol{\theta}^{(i)},\mathbf{d})-\ln q(\mathbf{y}^{(i)};\boldsymbol{\lambda})\right\}.$ (31)
@@ 4 Numerical Experiments
We demonstrate vOED-NFs across four design cases involving models of varying complexity: (1) a nonlinear model with 3 parameters and 1 design variable; (2) a linear model with 21 parameters and 400-dimensional design vector; (3) a PDE model for cathodic electrophoretic deposition; and (4) a stochastic model of aphid population with an implicit likelihood.
@@ 4.1 Case 1: Low-dimensional nonlinear design
Consider a nonlinear observation model [^1] There is a typo to (32) in the journal version of this paper. The equation provided here is the correct one and consistent with our results and code.
$y=G(\boldsymbol{\theta},d)+\epsilon\qquad\text{where}\qquad G(\boldsymbol{\theta},d)=\theta_{1}^{3}d^{2}+\theta_{2}e^{-|0.2-d|}+\sqrt{\theta_{3}^{2}d\times 2},$ (32)
@@
Figure 2(a) presents the EIG estimates across the discretized design space using (a) (Hi-NMC) a high-quality $\widehat{U}_{\text{NMC}}$ with $N_{\text{out}}=2\times 10^{4}$ and $N_{\text{in}}=2\times 10^{4}$ samples, equating to $4\times 10^{8}$ forward model evaluations total, (b) (Lo-NMC) a low-quality $\widehat{U}_{\text{NMC}}$ with $N_{\text{out}}=200$ and $N_{\text{in}}=200$ samples, equating to $4\times 10^{4}$ forward model evaluations total, (c) (Re-NMC) a NMC estimator with $N_{\text{out}}=N_{\text{in}}=2\times 10^{4}$, where the inner loop reuses the outer loop samples, equating to $2\times 10^{4}$ forward model evaluations total, (d) (vOED-G) EIG lower bound $\widehat{U}_{L}$ using Gaussian distributions optimized with $N_{\text{opt}}=2\times 10^{4}$ samples, and (e) (vOED-NFs) EIG lower bound $\widehat{U}_{L}$ using NFs optimized with $N_{\text{opt}}=2\times 10^{4}$ samples. Furthermore, the lower bound estimate $\widehat{U}_{L}$ at each $d$ is evaluated using $N=1\times 10^{4}$ samples. For (d) and (e), the total number of forward model evaluation is therefore $2\times 10^{4}+1\times 10^{4}=3\times 10^{4}$. vOED-NFs adopts $T=5$ sets of complete transformations of the cINN described in Sec. 3.1. Additional training details and hyperparameter choices for (d) and (e) can be found in C.1.
In Fig. 2(a), Hi-NMC is designated as the reference solution. In comparison to this reference, Lo-NMC exhibits moderate bias while vOED-NFs performs extremely well. Re-NMC yields near-identical results as Hi-NMC; while faster, it still requires an explicit likelihood and does not provide approximate posterior distributions, just like the other NMC methods. vOED-G, due to its much simplistic Gaussian variational distributions, deviates significantly from the expected utility trend and misidentifies the optimal design. However, the training cost for vOED-NFs is also higher than vOED-G, as illustrated through the training convergence history for $d=1.0$ in Fig. 2(b). Here vOED-G shows a very rapid convergence to $\boldsymbol{\lambda}^{*}$, while vOED-NFs takes longer to stabilize and experiences greater fluctuation. This timing difference induced by optimization is negligible when the overall computations are dominated by the forward model runs, but can become significant when the forward model is inexpensive. Additional information regarding the time required for EIG estimation is given in G.
Figure 2: Case 1. EIG estimates and sample training convergence plot. The Re-NMC plot line is right on top of the High-NMC plot line, as they produce near-identical results.
@@
Figure 5 investigates the impact of the number of transformations and optimization sample size $N_{\text{opt}}$ in vOED-NFs. The plots show the EIG lower bound estimates $\widehat{U}_{L}$ when adopting $T=\{1,3,5,7\}$ sets of complete cINN transformations, and using $N_{\text{opt}}=\{5\times 10^{3},10^{4},5\times 10^{4}\}$. The EIG lower bound estimates are still evaluated with $N=10^{4}$ samples. In general, employing 3 transformations is adequate to identify the optimal design $d^{\ast}=1$. However, when $N_{\text{opt}}$ is small, adding more transformations does not provide significant improvement. When $N_{\text{opt}}$ is increased, the value from adding transformations is more notable and stabilized around $T=5$; this is consistent with previous work [55] that reported composing more than 5 transformations provides only slight additional benefits. This diminishing return is likely due to the much larger number of NFs parameters when a high number of transformations is used, making them more prone to overfitting. Additional testing and discussions for the $N_{\text{opt}}=5\times 10^{3}$ case can be found in C.1.
@@ 4.2 Case 2: High-dimensional linear design
where $y_{j}\in\mathbb{R}$ is the $j$th observation, $\mathbf{d}_{j}\in\mathbb{R}^{p}$ is its corresponding design vector, $\epsilon_{j}\sim\mathscr{N}(0,\sigma^{2})$ is a Gaussian noise random variable, and $\boldsymbol{\theta}=\{\mathbf{w},\sigma\}$ encompasses the $p$-dimensional regression coefficient vector $\mathbf{w}$ and noise standard deviation $\sigma$. We set $p=20$ and $n=20$, hence the parameter dimension is 20 and the total design dimension is 400. Independent priors are adopted for $\mathbf{w}_{j}\sim\text{Laplace}(0,1)$ and $\sigma\sim\text{Exp}(1)$. Lastly, a design constraint $||\mathbf{d}_{j}||_{1}=1$ is imposed to reflect a resource budget; this is implemented through re-normalization of $\mathbf{d}$ after each SGA update in order to map it back to the feasible region. During the joint optimization of $\mathbf{d}$ and $\boldsymbol{\lambda}$, the total number of forward model runs (i.e., the total number of $(\boldsymbol{\theta}^{(i)}$, $\mathbf{y}^{(i)})$ pairs across all design epochs) is fixed to $10^{6}$. This is the same computational budget used by [15] that adopted Gaussian and Gamma variational distributions for $\mathbf{w}$ and $\sigma$, respectively; we refer to the setup in [15] as vOED-GG. Further details on the computational setup can be found in D.1.
Optimization results are presented in Table 1, with the middle column showing the mean $\pm$ one standard deviation of 10 lower bound estimates $\widehat{U}_{L}$ (each evaluated with $N=10^{4}$ samples) at the optimal design found by vOED-GG and vOED-NFs, and the right column showing the mean $\pm$ one standard deviation of 10 high-quality NMC estimates (each using $N_{\text{out}}=10^{4}$ and $N_{\text{in}}=10^{4}$) at each method’s optimal design to provide an accurate EIG estimate and using a common estimation method. We observe that vOED-NFs’s optimal design achieves a tighter lower bound estimate and also a higher NMC estimate, indicating that a better design has been found than vOED-GG. However, the vOED-NFs lower bound estimate is not tight, and the remaining discrepancy between $\widehat{U}_{L}$ and $\widehat{U}_{\text{NMC}}$ may be contributed by a number of factors: (1) the functional form of cINN; (2) the number of cINN transformations; (3) the functional representation of $s$’s and $t$’s; (4) the possibility to converge to a local optimum during optimization; and (5) the finite sample sizes.
@@
vOED-GG   12.07 (0.11)   24.28 (0.24)
vOED-NFs   20.94 (0.19)   24.91 (0.24)
Table 1: Case 2. Middle column shows the mean of 10 lower bound estimates $\widehat{U}_{L}$ (each is evaluated with $N=10000$ samples) at the optimal design found by vOED-GG and vOED-NFs, and the right column shows the mean of 10 high-quality NMC estimates (each with $N_{\text{out}}=10000$ and $N_{\text{in}}=10000$) at the optimal designs. The ()($\cdot$)( ⋅ ) values are one standard deviation estimates for the estimators.
In order to compare the posterior approximation capabilities of the two approaches, we look at their approximate posteriors at a common design with each $\mathbf{d}_{j}=\mathbf{e}_{j}$ set to the standard unit vector in the $j$th dimension, and at three $\mathbf{y}$ samples. Figures 6 and 7 present the marginal posteriors for $\{w_{1},w_{2},w_{3},w_{4},\sigma\}$ using No-U-Turn Hamiltonian Monte Carlo (NUTS-HMC) [61] as the reference posterior distribution, along with the approximations obtained from vOED-GG and vOED-NFs. From the figure, it is evident that the vOED-NFs posteriors can capture the non-Gaussian structures much better than vOED-GG’s Gaussian/Gamma distributions, which is confirmed by the empirical KL divergence from the NFs posteriors to MCMC posteriors.
Figure 6: Case 2. Comparison of marginal posteriors obtained from NUTS-HMC and vOED-GG [24] at design where $\mathbf{d}_{j}=\mathbf{e}_{j}$, with KLD denotes the empirical KL divergence from NFs posteriors to MCMC posteriors.
Figure 7: Case 2. Comparison of marginal posteriors obtained from NUTS-HMC and vOED-NFs [24] at design where $\mathbf{d}_{j}=\mathbf{e}_{j}$, with KLD denotes the empirical KL divergence from NFs posteriors to MCMC posteriors.